# A Survey on Inference Optimization Techniques for Mixture of Experts Models

### v1

@@ Abstract.
The emergence of large-scale Mixture of Experts (MoE) models has marked a significant advancement in artificial intelligence, offering enhanced model capacity and computational efficiency through conditional computation. However, the deployment and inference of these models present substantial challenges in terms of computational resources, latency, and energy efficiency. This comprehensive survey systematically analyzes the current landscape of inference optimization techniques for MoE models across the entire system stack. We first establish a taxonomical framework that categorizes optimization approaches into model-level, system-level, and hardware-level optimizations. At the model level, we examine architectural innovations including efficient expert design, attention mechanisms, various compression techniques such as pruning, quantization, and knowledge distillation, as well as algorithm improvement including dynamic routing strategies and expert merging methods. At the system level, we investigate distributed computing approaches, load balancing mechanisms, and efficient scheduling algorithms that enable scalable deployment. Furthermore, we delve into hardware-specific optimizations and co-design strategies that maximize throughput and energy efficiency. This survey not only provides a structured overview of existing solutions but also identifies key challenges and promising research directions in MoE inference optimization. Our comprehensive analysis serves as a valuable resource for researchers and practitioners working on large-scale deployment of MoE models in resource-constrained environments. To facilitate ongoing updates and the sharing of cutting-edge advances in MoE inference optimization research, we have established a repository accessible at https://github.com/MoE-Inf/awesome-moe-inference/.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Mixture of Experts (MoE) has emerged as a promising architectural solution to address scaling challenges in machine learning (shazeer2017outrageously, ). Originally introduced by Jacobs et al.(articleJacobs, ) in the early 1990s as a method for learning subtasks in neural networks, Numerous MoE-based models (MetaPi, ; NIPS20009fdb62f9, ; Eigen2013LearningFR, ) have been developed over the years. In the era of large language models, MoE has again experienced a renaissance (jiang2024mixtral, ; abdin2024phi, ; dai2024deepseekmoe, ; sun2024hunyuan, ). The core principle of MoE is to distribute the model’s capacity across multiple specialized sub-networks, or experts, with a learned gating mechanism that selectively activates only the relevant experts for each input. This approach allows models to maintain a large parameter count while keeping computational costs manageable through sparse activation. Recent implementations, such as Mixtral 8x7B (jiang2024mixtral, ), Switch Transformers (fedus2022switch, ) and GShard (lepikhin2020gshard, ), have demonstrated the effectiveness of this strategy in scaling language models to trillions of parameters while maintaining reasonable computational requirements.
{forest}
[…]
Figure 1. Taxonomy of MoE inference optimization literature.
@@
Table 1. A List of SOTA MoEs. Attribute Param. represents the number of total parameters. Experts according to the format of number of Activation Experts/Total Experts/Shared Experts. #L represents the number of hidden layers, #H represents the number of attention heads. $d_{model}$ is the hidden size, $d_{ffn}$ is the intermediate size of FFNs, $d_{expert}$ is the intermediate size of FFN experts.
@@
The success of MoE in scaling models has led to its adoption in various state-of-the-art systems. For example, Google’s GLaM (du2022glam, ) outperforms GPT-3 while using significantly fewer computational resources during inference. Similarly, Mixtral 8x7B (jiang2024mixtral, ), a recent open-source MoE model, has demonstrated competitive performance compared to much larger dense models, while maintaining efficient inference characteristics. Table 1 summarizes recent state-of-the-art open-source MoE models that have garnered significant attention, further highlighting the strong potential of the MoE architecture. These successes have sparked widespread interest in MoE across both academia and industry, leading to innovations in model design (chowdhury2023patch, ; wang2020deep, ; zhang2022mixture, ), training techniques (liu2023moelora, ; dou2024loramoe, ; gao2024higher, ), and deployment strategies (chen2022ta, ; cao2024moelight, ; 10528887, ).
@@
However, the efficient deployment of MoE models for inference presents unique and significant challenges (tang2024hobbit, ; yi2023edgemoe, ; zhong2024adapmoe, ; hwang2024pre, ). The dynamic nature of expert activation patterns introduces complexity in resource management and scheduling that is not present in traditional dense models. These challenges span multiple levels: at the model level, the design of efficient expert architectures and routing mechanisms directly impacts inference performance; at the system level, managing distributed computation and load balancing becomes increasingly complex; and at the hardware level, specialized acceleration techniques are needed to handle the sparse computation patterns.
Numerous methods have been developed to address these challenges in MoE deployment and inference (jin2024moh, ; xie2024moe, ; DeepSpeedmoe, ; Sarkar2024edgemoe, ). While the rapid growth of research in this field demonstrates its importance, it can also make it difficult to identify key trends and best practices. A critical gap in existing literature is the absence of a systematic framework for analyzing and developing comprehensive inference optimization solutions for MoE models.
@@
To bridge this gap, our work offers a comprehensive survey of inference optimization techniques for MoE models. We propose a taxonomical framework that categorizes optimization approaches into model-level, system-level, and hardware-level optimizations, as illustrated in Figure 1. This framework provides a structured approach to understanding and comparing different optimization techniques. While there are related surveys on LLM efficiency (yuan2024llm, ; li2024llm, ; li2024large, ; treviso2023efficient, ; wan2023efficient, ; bai2024beyond, ; zhou2024survey, ; xu2024survey, ) and MoE architectures (202408.0583, ; cai2024survey, ; fedus2022review, ), our work is the first to specifically focus on inference optimization techniques for MoE models. We systematically analyze optimization approaches at different abstraction levels, from model architecture to hardware acceleration, providing a valuable resource for researchers and practitioners working deploy MoE models for different real-world applications.
@@ 2. Fundamentals of Mixture of Experts
Mixture of Experts (MoE) represents a significant architectural paradigm in neural networks, particularly in large language models, where it enables conditional computation through sparse activation mechanisms (cai2024survey, ). At its core, an MoE architecture consists of a routing network $R(x)$ and a set of expert networks ${E_{1},E_{2},...,E_{N}}$, where $N$ denotes the total number of experts. The fundamental principle of MoE can be expressed as $y=\sum_{i=1}^{N}g_{i}(x)\cdot E_{i}(x)$, where $g_{i}(x)$ represents the gating function for expert $i$, and $E_{i}(x)$ is the output of expert $i$.
@@
In the inference process of a sparse MoE model, computation follows a three-stage pipeline. First, the router computes expert selection probabilities:
@@ 3. Model-level Optimizations
Model-level optimizations aim to enhance the inherent structure and efficiency of MoE models through systematic improvements in architecture, parameter optimization, and algorithmic design. These optimizations can be broadly categorized into three main areas: efficient model architecture design, model compression techniques, and algorithmic improvements. Architecture design focuses on developing more efficient expert and attention structures, while compression techniques aim to reduce model size and memory footprint through methods such as pruning, quantization, and knowledge distillation. Algorithmic improvements concentrate on enhancing the dynamic aspects of MoE models, including routing mechanisms and expert combination strategies. Figure 3 illustrates the detailed structure of this section.
@@ 3.1.1. MoE-based Attention Design
In addition to the typical application of the MoE structure in the FFN module of transformer layer, current studies explore how to incorporate MoE into the Attention module for improved performance. MAE (peng-etal-2020-mixture, ) was the first to explain the multi-head attention mechanism from the MoE perspective, using a learned gating function to activate different experts for different inputs, with each expert consisting of $n-1$ heads. To further optimize MoE-based attention modules, existing research proposes various structures. MoA (zhang2022mixture, ) and BAM (zhang2024bam, ) select $k$ heads for a given input and share key projection and value projection weights among all heads, while SwitchHead (csordas2023switchhead, ) shares key projection and query projection weights to enhance computational efficiency. MoH (jin2024moh, ) introduces shared heads and a two-stage routing process to further improve the standard MoE method, offering an advantage over MoA. Building upon MoA, ModuleFormer (shen2023moduleformer, ) extends sparse modules to both the attention and feed-forward layers, allowing for the easy addition and removal of modules. Inspired by MoA and ModuleFormer, JetMoE-8B (shen2024jetmoe, ) develops a powerful open-source model featuring sparse attention and sparse feed-forward layers, while DS-MoE (pan2024dense, ) proposes a hybrid dense training and sparse inference framework for efficient training and inference. Additionally, SUT (tan2023sparse, ) and MoEUT (csordas2024moeut, ) use sparse attention and sparse feed-forward layers to construct the efficient Sparse Universal Transformer model, which shares parameters across all layers.
@@ 3.1.2. MoE-based FFN Design
To enhance the efficiency of MoE-based models, current research explores various variants of the standard MoE module. MoE++ (jin2024moe++, ) introduces three types of zero-computation experts based on standard experts, aimed at reducing computational overhead. SCoMoE (xiongscomoe, ) leverages a structured all-to-all communication approach, inspired by hierarchical communication topologies, to reduce communication costs during parallel MoE computation. Pre-gated MoE (hwang2024pre, ) proposes a pre-gate MoE module that prefetches the required experts to improve inference speed on memory-constrained devices. COMET (ibrahim2023comet, ) introduces a tree-based sparse expert selection mechanism to optimize the traditional gating module, which typically relies on a linear approach. Additionally, MoELoRA (luo2024moelora, ) reimagines LoRA as a MoE for more parameter-efficient fine-tuning.
@@ 3.2. Model Compression Techniques
Figure 4. Compression techniques in MoE models.
@@ 3.2.1. Expert Pruning
Due to the large number of parameters in MoE-based models, current research explores pruning methods to reduce the number of parameters in MoE experts. These methods are generally divided into structured and unstructured pruning. Most studies focus on structured expert pruning, which aims to reduce the number of experts in the MoE layer while maintaining model accuracy. As shown in Figure 4-(a), some approaches directly remove unimportant experts (left side), while others merge groups of experts into a single expert (right side). TSEP (chen2022task, ) removes non-professional experts for the target downstream task, while fine-tuning professional experts to preserve model performance. NAEE (lu2024not, ) eliminates unimportant experts by evaluating expert combinations on a small calibration dataset to minimize accuracy loss, while UNCURL (sarkar2024revisiting, ) reduces the number of experts based on MoE router logits. PEMPE (chowdhury2024provably, ) prunes experts that have smaller changes in the router’s $l_{2}$ norm between the pre-trained and fine-tuned models, while SEER-MoE (muzio2024seer, ) employs a heavy-hitters counting method for expert pruning. And MoE-$I^{2}$ (yang2024moe, ) proposes the Layer-wise Genetic Search and Block-wise KT-Reception Field with the non-uniform pruning ratio to prune unimportant experts. In addition to direct pruning, some studies utilize expert merging to reduce the number of experts. DEK (zhang2024diversifying, ) identifies and groups similar experts in feature space, then merges experts within the same group in weight space to reduce the expert count. EEP (liu2024efficient, ) introduces a two-stage approach to both prune and merge experts, reducing the total number of experts (saving GPU memory) and the number of active experts (accelerating inference). And MC-SMoE (li2023merge, ) divides experts into different groups based on routing policies and then merges each group into one expert. For unstructured pruning, MoE-Pruner (xie2024moe, ) prunes weights with the smallest magnitudes, weighted by the corresponding input activations and router weights. Moreover, STUN (lee2024stun, ) combines structured and unstructured pruning to achieve better performance than unstructured pruning alone. MoE-Compression (he2024demystifying, ) proposes a unified framework for MoE compression, using both structured and unstructured pruning methods to achieve significant inference speedup with minimal accuracy loss.

### v2

@@ Abstract.
The emergence of large-scale Mixture of Experts (MoE) models represents a significant advancement in artificial intelligence, offering enhanced model capacity and computational efficiency through conditional computation. However, deploying and running inference on these models presents significant challenges in computational resources, latency, and energy efficiency. This comprehensive survey analyzes optimization techniques for MoE models across the entire system stack. We first establish a taxonomical framework that categorizes optimization approaches into model-level, system-level, and hardware-level optimizations. At the model level, we examine architectural innovations including efficient expert design, attention mechanisms, various compression techniques such as pruning, quantization, and knowledge distillation, as well as algorithm improvement including dynamic routing strategies and expert merging methods. At the system level, we investigate distributed computing approaches, load balancing mechanisms, and efficient scheduling algorithms that enable scalable deployment. Furthermore, we delve into hardware-specific optimizations and co-design strategies that maximize throughput and energy efficiency. This survey provides both a structured overview of existing solutions and identifies key challenges and promising research directions in MoE inference optimization. To facilitate ongoing updates and the sharing of cutting-edge advances in MoE inference optimization research, we have established a repository accessible at https://github.com/MoE-Inf/awesome-moe-inference/.
Mixture of Experts, Large Language Models, Inference Optimization
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Neural networksccs: Computer systems organization Neural networksccs: Computer systems organization Embedded and cyber-physical systems
@@ 1. Introduction
Mixture of Experts (MoE) has emerged as a promising architectural solution to address scaling challenges in machine learning (shazeer2017outrageously, ). Originally introduced by Jacobs et al.(articleJacobs, ) in the early 1990s as a method for learning subtasks in neural networks. Numerous MoE-based models (MetaPi, ; NIPS20009fdb62f9, ; Eigen2013LearningFR, ) have been developed over the years. In the era of large language models, MoE has again experienced a renaissance (jiang2024mixtral, ; abdin2024phi, ; dai2024deepseekmoe, ; sun2024hunyuan, ). The core principle of MoE is to distribute the model’s capacity across multiple specialized sub-networks, or experts, with a learned routing mechanism that selectively activates only the relevant experts for each input. This approach allows models to maintain a large parameter count while keeping computational costs manageable through sparse activation. Recent implementations, such as Mixtral 8x7B (jiang2024mixtral, ), DeepSeek-V3 (liu2024deepseek, ) and DBRX (DBRX, ), have demonstrated the effectiveness of this strategy in scaling language models to trillions of parameters while maintaining reasonable computational requirements.
@@
DeepSeek-V3 (liu2024deepseek, ) 671B 8/256/1 61 128 7168 18432 2048 DeepSeek-AI 2024.12
MiniMax-Text-01 (minimax2025minimax01scalingfoundationmodels, ) 456B 2/32/0 80 64 6144 9216 9216 MiniMax-AI 2025.1
@@
Table 1. A List of SoTA MoEs. Param. represents the number of total parameters. Experts are listed according to the format of the number of activation experts, total experts, and shared experts. #L represents the number of hidden layers, #H represents the number of attention heads. $d_{model}$ is the hidden size, $d_{ffn}$ is the intermediate size of FFNs, $d_{expert}$ is the intermediate size of FFN experts.
{forest}
[…]
Figure 1. Taxonomy of MoE inference optimization.
@@
The success of MoE in scaling models has led to its adoption in various state-of-the-art systems. For example, Google’s GLaM (du2022glam, ) outperforms GPT-3 while using significantly fewer computational resources during inference. Similarly, Mixtral 8x7B (jiang2024mixtral, ) has demonstrated competitive performance compared to much larger dense models, while maintaining efficient inference characteristics. DeepSeek-V3 (liu2024deepseek, ), a recent open-source MoE model, has surpassed other open-source alternatives and demonstrated performance comparable to prominent closed-source models such as GPT4-o and Claude-3.5-Sonnet. Table 1 summarizes recent state-of-the-art open-source MoE models that have garnered significant attention, showcasing their rapid evolution and widespread adoption across major tech companies and research institutions. The models range in size from 6.7B to 671B parameters, with varying architectures characterized by different numbers of experts, hidden layers, and attention heads. The chronological progression from 2022 to late 2024 demonstrates increasing model sizes and architectural sophistication, with recent models like DeepSeek-V3 pushing the boundaries in terms of parameter count and model performance. This trend further highlights the strong potential of the MoE architecture in advancing the field of large language models. These successes have sparked widespread interest in MoE across both academia and industry, leading to innovations in model design (chowdhury2023patch, ; wang2020deep, ; zhang2022mixture, ), training techniques (liu2023moelora, ; dou2024loramoe, ; gao2024higher, ), and deployment strategies (chen2022ta, ; cao2024moelight, ; 10528887, ).
@@
However, the efficient deployment of MoE models for inference presents unique and significant challenges (tang2024hobbit, ; yi2023edgemoe, ; zhong2024adapmoe, ; hwang2024pre, ). The dynamic nature of expert activation patterns introduces complexity in resource management and scheduling that is not present in traditional dense models. At the model level, the design of efficient expert architectures faces challenges in balancing model capacity with computational efficiency, while routing mechanisms must optimize expert selection and load distribution. The system-level challenges are particularly complex: distributed computation requires sophisticated scheduling algorithms to manage expert placement and activation, load balancing must handle dynamic workload variations across experts, and memory management needs to efficiently handle the loading and unloading of expert parameters. Hardware-level challenges stem from the fundamental mismatch between traditional hardware architectures optimized for dense computation and the sparse, dynamic nature of MoE inference. This necessitates specialized acceleration techniques to handle sparse computation patterns, manage memory bandwidth efficiently, and provide flexible computation capabilities for dynamic expert switching. Communication overhead in distributed settings presents another significant challenge, particularly when experts are distributed across different devices or nodes, requiring careful optimization of data movement and synchronization.
@@
Numerous methods have been developed to address these challenges in MoE deployment and inference (jin2024moh, ; xie2024moe, ; DeepSpeedmoe, ; Sarkar2024edgemoe, ). While the rapid growth of research in this field demonstrates its importance, it can also make it difficult to identify key trends and best practices. A critical gap in the existing literature is the absence of a systematic framework for analyzing and developing comprehensive inference optimization solutions for MoE models. To bridge this gap, our work offers a comprehensive survey of inference optimization techniques for MoE models. We propose a taxonomical framework that categorizes optimization approaches into model-level, system-level, and hardware-level optimizations, as illustrated in Figure 1. This framework provides a structured approach to understanding and comparing different optimization techniques. While there are related surveys on LLM efficiency (yuan2024llm, ; li2024llm, ; li2024large, ; treviso2023efficient, ; wan2023efficient, ; bai2024beyond, ; zhou2024survey, ; xu2024survey, ) and MoE architectures (202408.0583, ; cai2024survey, ; fedus2022review, ), our work is the first to specifically focus on inference optimization techniques for MoE models. We systematically analyze optimization approaches at different abstraction levels, from model architecture to hardware acceleration, providing a valuable resource for researchers and practitioners working deploy MoE models for different real-world applications.
@@ 2. Fundamentals of Mixture of Experts
MoE represents a significant architectural paradigm in neural networks, particularly in large language models, where it enables conditional computation through sparse activation mechanisms (cai2024survey, ). At its core, an MoE architecture consists of a routing network $R(x)$ and a set of expert networks ${E_{1},E_{2},...,E_{N}}$, where $N$ denotes the total number of experts. The fundamental principle of MoE can be expressed as $y=\sum_{i=1}^{N}g_{i}(x)\cdot E_{i}(x)$, where $g_{i}(x)$ represents the gating function for expert $i$, and $E_{i}(x)$ is the output of expert $i$.
@@
The typical MoE model inference procedure follows a sequence of expert selection, parallel computation, and output aggregation. First, the router computes expert selection probabilities:
@@
The MoE architecture offers several compelling advantages that have contributed to its growing adoption in modern neural networks. First, the conditional computation mechanism enables significant computational savings compared to dense models of similar capacity. By activating only a subset of experts for each input token, MoE models can process information more efficiently while maintaining high performance levels. This is particularly valuable in resource-constrained environments or when scaling to larger model sizes.
Second, the specialization of individual experts allows for more nuanced and accurate processing of different input patterns. Each expert can develop specialized knowledge for specific aspects of the input space, leading to more refined and targeted computations. This specialization is particularly beneficial in language models, where different experts can focus on various linguistic patterns, domains, or tasks.
Third, the dynamic routing mechanism enables adaptive computation based on input complexity. More challenging or nuanced inputs can engage multiple experts with complementary specializations, while simpler inputs might require only a small subset of experts. This adaptive resource allocation helps optimize the computation-performance trade-off and potentially improves both efficiency and effectiveness.
@@ 3. Model-level Optimizations
Model-level optimizations aim to enhance the inherent structure and efficiency of MoE models through systematic improvements in architecture, parameter optimization, and algorithmic design. These optimizations can be broadly categorized into three main areas: efficient model architecture design, model compression techniques, and algorithmic improvements. Architecture design focuses on developing more efficient expert and attention structures, while compression techniques aim to reduce model size and memory footprint through methods such as pruning, quantization, and knowledge distillation. Algorithmic improvements concentrate on enhancing the dynamic aspects of MoE models, including routing mechanisms and expert combination strategies. Figure 3 illustrates the detailed taxonomy of model-level optimization that is described in this section.
@@ 3.1.1. MoE-based Attention Design
In addition to the typical application of the MoE structure in the FFN module of the transformer layer, current studies explore how to incorporate MoE into the Attention module for improved performance. MAE (peng-etal-2020-mixture, ) was the first to explain the multi-head attention mechanism from the MoE perspective, using a learned gating function to activate different experts for different inputs, with each expert consisting of $n-1$ heads. To further optimize MoE-based attention modules, existing research proposes various structures. MoA (zhang2022mixture, ) and BAM (zhang2024bam, ) select $k$ heads for a given input and share key projection and value projection weights among all heads, while SwitchHead (csordas2023switchhead, ) shares key projection and query projection weights to enhance computational efficiency. MoH (jin2024moh, ) introduces shared heads and a two-stage routing process to further improve the standard MoE method, offering an advantage over MoA. Building upon MoA, ModuleFormer (shen2023moduleformer, ) extends sparse modules to both the attention and feed-forward layers, allowing for the easy addition and removal of modules. Inspired by MoA and ModuleFormer, JetMoE-8B (shen2024jetmoe, ) develops a powerful open-source model featuring sparse attention and sparse feed-forward layers, while DS-MoE (pan2024dense, ) proposes a hybrid dense training and sparse inference framework for efficient training and inference. Additionally, SUT (tan2023sparse, ) and MoEUT (csordas2024moeut, ) use sparse attention and sparse feed-forward layers to construct the efficient Sparse Universal Transformer model, which shares parameters across all layers.
@@ 3.1.2. MoE-based FFN Design
To enhance the efficiency of MoE-based models, current research explores various variants of the standard MoE module. MoE++ (jin2024moe++, ) introduces three types of zero-computation experts based on standard experts, aimed at reducing computational overhead. SCoMoE (xiongscomoe, ) leverages a structured all-to-all communication approach, inspired by hierarchical communication topologies, to reduce communication costs during parallel MoE computation. Pre-gated MoE (hwang2024pre, ) proposes a pre-gated MoE module that prefetches the required experts to improve inference speed on memory-constrained devices. COMET (ibrahim2023comet, ) introduces a tree-based sparse expert selection mechanism to optimize the traditional gating module, which typically relies on a linear approach. Additionally, MoELoRA (luo2024moelora, ) reimagines LoRA as a MoE for more parameter-efficient fine-tuning.
@@ 3.2. Model Compression Techniques
Figure 4. Compression techniques for MoE models.
@@ 3.2.1. Expert Pruning
Sparsity TSA Datasets of Fine-tuning Structured Unstructured
Delete Merge
TSEP (chen2022task, ) 96.875% S GLUE (wang2018glue, ) SQuAD (rajpurkar2016squad, )
NAEE (lu2024not, ) 50% S MetaMathQA (yu2023metamath, )
UNCURL (sarkar2024revisiting, ) 75% S FLAN (longpre2023flan, )
PEMPE (chowdhury2024provably, ) 75% A CIFAR-10 CIFAR-100 (krizhevsky2009learning, ) ImageNet (russakovsky2015imagenet, )
SEER-MoE (muzio2024seer, ) 25% S MMLU (hendrycks2020measuring, ) SST5 (socher2013recursive, )
MoE-$I^{2}$ (yang2024moe, ) 53.98% A Alpaca (taori2023stanford, )
DEK (zhang2024diversifying, ) 75% S C4 (raffel2020exploring, )
EEP (liu2024efficient, ) 75% A NA
MC-SMoE (li2023merge, ) 75% S eight datasets
MoE-Pruner (xie2024moe, ) 50% A C4 (raffel2020exploring, )
STUN (lee2024stun, ) 70% A C4 (raffel2020exploring, )
MoE-compression (he2024demystifying, ) 50% A Lima (zhou2024lima, ) MetaMathQA (yu2023metamath, )
Table 2. A comparison of pruning methods. Sparsity indicates the max ratio of pruning. TSA represents task specific or agnostic. EEP (liu2024efficient, ) does not require additional fine-tuning. Finetuning datasets of MS-SMoE (li2023merge, ) are SST2 (socher2013recursive, ) MRPC (dolan2005automatically, ) MultiRC (khashabi2018looking, ) COPA (roemmele2011choice, ) WinoGrande (sakaguchi2021winogrande, ) SQuAD (rajpurkar2016squad, ) WikiQA (yang2015wikiqa, )and HotpotQA (yang2018hotpotqa, ).
@@
Due to the large number of parameters in MoE-based models, current research explores pruning methods to reduce the number of parameters in MoE experts. These methods are generally divided into structured and unstructured pruning. Most studies focus on structured expert pruning, which aims to reduce the number of experts in the MoE layer while maintaining model accuracy. As shown in Figure 4-(a), some approaches directly remove unimportant experts (left side), while others merge groups of experts into a single expert (right side). TSEP (chen2022task, ) removes non-professional experts for the target downstream task while fine-tuning professional experts to preserve model performance. They also demonstrate the superiority of the eager expert pruning paradigm over other possible solutions like two-pass optimization or staged expert pruning. NAEE (lu2024not, ) eliminates unimportant experts by evaluating expert combinations on a small calibration dataset to minimize accuracy loss, simultaneously reduce model sizes and increase inference speed. while UNCURL (sarkar2024revisiting, ) reduces the number of experts based on MoE router logits. Some insights useful to model design choices considering task-specific inference optimization can give guidance for later stages. PEMPE (chowdhury2024provably, ) prunes experts that have smaller changes in the router’s $l_{2}$ norm between the pre-trained and fine-tuned models, and will conduct experiments with other compression techniques on SOTA vision MoEs. SEER-MoE (muzio2024seer, ) employs a heavy-hitters counting method for expert pruning, then with regularization-based fine-tuning reaches further expert pruning. MoE-$I^{2}$ (yang2024moe, ) proposes the Layer-wise Genetic Search and Block-wise KT-Reception Field with the non-uniform pruning ratio to prune unimportant experts.
In addition to direct pruning, some studies utilize expert merging to reduce the number of experts. For unstructured pruning, MoE-Pruner (xie2024moe, ) prunes weights with the smallest magnitudes, weighted by the corresponding input activations and router weights. The pruned MoE models can benefit from a pre-trained teacher model through expert-wise knowledge distillation and have compatibility with structured pruning. Moreover, STUN (lee2024stun, ) combines structured and unstructured pruning to achieve better performance than unstructured pruning alone, utilizing the sparse character of MoEs based on behavior similarity in a greedy manner. MoE-Compression (he2024demystifying, ) proposes a unified framework for MoE compression, using both structured and unstructured pruning methods to achieve significant inference speedup with minimal accuracy loss.
Another potential approach is merging outdated experts based on their parameters to achieve better performance. Branch-Train-Merge (li2022branch, ) independently trains different parts of the model on distinct subsets of data, eliminating the need for large-scale multi-node synchronization typically required in traditional LLM training. Building on this, Branch-Train-Mix (sukhbaatar2024branch, ) trains multiple copies of a seed LLM to specialize in multiple domains in an asynchronous and parallel fashion, then merges the parameters of the MoE layer to create a unified model that can be further trained. The second finetuning stage makes the final LLM more performant. They also find that their approach is more computing efficient compared to the dense model or specialized MoE model on training. Park et al. (park2024learning, ) observed that introducing a shared layer in the MoE could lead to performance degradation. To address this, they trained merged experts that learned the same features in different ways, improving their ability to generalize and mitigating catastrophic forgetting during incremental learning of multi-domain tasks. MC-SMoE (li2023merge, ) divides experts into different groups based on routing policies and then merges each group into one expert. HC-SMoE (chen2024retraining, ) is a framework that uses hierarchical clustering to merge experts without requiring retraining and can be applied in a task-agnostic manner. During inference, MEO (he2023merging, ) systematically investigates the computational cost of MoE. In the drop-in replacement algorithm, they first merge the selected expert parameters and then perform efficient computation. FoE (wang2023fusing, ) fuses the outputs of expert models by leveraging their complementary knowledge of the data distribution, framing it as a supervised learning instance. DEK (zhang2024diversifying, ) identifies and groups similar experts in feature space, then merges experts within the same group in weight space to reduce the expert count. EEP (liu2024efficient, ) introduces a two-stage approach to both prune and merge experts, reducing the total number of experts (saving GPU memory) and the number of active experts (accelerating inference). Inspired by the concept of knowledge transfer in multi-task learning, HyperMoE (zhao-etal-2024-hypermoe, ) proposes a novel expert network that further increases sparsity while utilizing information from unselected experts as supplementary input. In practice, they capture the contextual information of experts to compensate for the performance loss of transferring knowledge to specific experts. LiteMoE (zhuang2024litemoe, ) retains the most critical experts based on the application’s characteristics, merges secondary experts, and obtains the final sparse model without retraining. This approach enables efficient deployment of lightweight submodels on resource-constrained mobile devices.